OpenStudy (anonymous):

Please i really need help this is hard. Which value is an extraneous solution to the equation?|x|=-5x+24 A. -6. B. 4 C. 6. D. 7

OpenStudy (anonymous):

|a| = b a = b or a = -b

OpenStudy (anonymous):

|x| = -5x + 24 x = -5x + 24 or x = 5x - 24 6x = 24 or 4x = 24 x = 4 or x = 6 Substitute the values back into the equation |4| = -5(4) + 24 4 = 4 |6| = -5(6) + 24 6 = -6 So, the extra answer is x = 6.